According to this table:

Microsoft Corporation (MSFT) NasdaqGS- NasdaqGS Real Time Price. Currency in USD Add to watchlist 72.40-1.25(-1.70%) 72.40 0.00 (0.00%) At close: 4:00PM EDT After hours: 4:20PM EDT Summary Conversations Statistics ProfileFinancials Options Holders Historical Data Analysts Previous Close Open Bid Ask Days Range 52 Week Range Volume 557.642B 1D 5D 1M 6M YTD 1Y 5Y Max Full screen 78.00 72.40 69.33 73.65 Market Cap 1.44 31.95 2.27 Oct 30, 2017 1.56 (2.12%) 2017-08-15 73.58 Beta 72.35 x 200 PERatio (TTM) 72.44 x 1200 EPS (TTM) 72.40 - 73.87 arnings Date 55.61 74.42 Dividend & Yield 60.67 52.00 22,203,456Ex-Dividend Date Aug 29, 16 Feb 20, 17 Avg. Volume 24,581,487 y Target Est 80.19

Question 7

The stock's beta of 1.44 tells you that its price is:

More volatile than the average stock
Projected to increase annually by 44%
Is higher than its dividend yield
44% higher than similar companies
Larger than alpha

Question 8

If the price of a share of Microsoft is $72.40 and the stock's P/E ratio is 31.95, what is the stock's earnings per share?

$1.44
$2.27
$3.19
$5.02
$7.24

Question 9

How much of an annual dividend does Microsoft pay per share?

$0.56
$1.25
$1.38
$1.56
$2.12

Question 10

If a share of Microsoft was selling for $80 and its dividend was $3.20 per share, what would its dividend yield be:

2%
3%
4%
5%
6%
Answer #1

Answer for Question 7

Beta tells us the movement of stock price with respect to market portfolio ie index Now it is 1.44 it means if market moves u by 1% then Price of stock will go up by 1.44% which tell us this stock is more volatile thana market

Hence option 1 is correct response

Answer for Question 8

P/E=MPS/EPS=31.95

EPS=MPS/31.95=72.4/31.95=$2.27

Option B

Ans for 9

Its already given in chart $1.56

Option D

Ans for 10

Dividend Yield=Dividend/Stock Price=3.2*100/80=4%

Option C